OpenStudy (anonymous):

Hmmm i think im messing up with every step...can someone please check my work? will post in a sec...

OpenStudy (anonymous):



OpenStudy (anonymous):

I keep getting decimals and if i simplify it further ill have:\[\frac{ 2(x+1)^{2} }{ 36.25 }\frac{ 2(y+1.5)^{2} }{ 36.25}=1\]

OpenStudy (anonymous):

@Hero @jim_thompson5910

OpenStudy (anonymous):

did I complete the square correctly?

OpenStudy (anonymous):

wow never mind the question is not asking for a standard formula lol its just asking about which conic this is, obviously a circle!!!